A denominator is 3 times the numerator of a fraction. If 4 is added to the denominator and 2 subtracted from the numerator, the resulting fraction is 3/4. Find the origina fraction.

Answer by checkley71(8403)

(x-2)/[(x+3)+4]=3/4
(x-2)/(x+7)=3/4 now cross multiply.
4(x-2)=3(x+7)
4x-8=3x+21
4x-3x=21+8
x=29 answer.
PROOF:
(29-2)/[(29+3)+4]=3/4
27/36=3/4
3/4=3/4